how do you find the center and radius for this equation for a circle
x^2+y^2-18x-18y+53=0

Answer by scott8148(6628)

complete the squares

(x^2 - 18x + 81) + (y^2 - 18y + 81) = -53 + 81 + 81

(x - 9)^2 + (y - 9)^2 = 109

center ___ (9,9)

radius ___ sqrt(109)